OpenStudy (liizzyliizz):

So in this question, Im not sure what i did wrong, for the answer I basically got the original equation written in slope intercept form. (Unless that is the correct answer? ) Write the equation of the line parallel to line 2x-3y=7 and containing the point (5,1).... Did i forget something?

jimthompson5910 (jim_thompson5910):

2x-3y=C 2(5)-3(1)=C 10-3=C 7 = C C = 7 So the equation parallel to 2x-3y=7 and that goes through (5,1) is 2x-3y=7. Make sure that there are no typos or other posting errors.

OpenStudy (liizzyliizz):

Nope no typos, I wonder if this was some sort of trick question, but yeah i got the same response.

jimthompson5910 (jim_thompson5910):

Yeah it's either a trick question or there's a typo.